OpenStudy (anonymous):

Find the exact value for arcsin(sin 630)

OpenStudy (anonymous):

find a point on the unit circle between \(-90\) and \(90\) that is coterminal with \(630\)

OpenStudy (p0sitr0n):

arcsin and sin are inverse function, thus they cancel out (since the existence of inverse implies bijectivity)

OpenStudy (anonymous):

cancel my foot!

OpenStudy (anonymous):

But the answer is supposed to be 201pi-603

OpenStudy (anonymous):

???

OpenStudy (anonymous):

are you working in degrees or radians?

OpenStudy (anonymous):

wait i mean 201pi-630

OpenStudy (anonymous):

in radians

OpenStudy (anonymous):

like hell

OpenStudy (zarkon):

"cancel my foot!" LOL

OpenStudy (anonymous):

when you write \[\arcsin(\sin(630))\] it is pretty much assumed you are working in degrees

OpenStudy (anonymous):

cancel \[\frac{8}{2}=4\] because the two's cancel \[x^2-2x+2x-4=x^2-4\] because the \(2x\) cancel \[\sqrt{x^2}=x\] because the square root and the square cancel \[e^{\ln(x)}=x\] because the \(e\) and the \(\ln(x)\) cancel what the heck

OpenStudy (zarkon):

Is that the word of the day :)

OpenStudy (anonymous):

\[\huge\cancel{\text{cancel}}\]

OpenStudy (anonymous):

so is the answer just 630?

OpenStudy (anonymous):

no it is \(-90\)

OpenStudy (anonymous):

\[630-360=270\\ 270-360=-90\]so \(-90\) is coterminal with \(630\) and that is the number you are looking for, the number between \(-90\) and \(90\) coterminal with \(630\)

OpenStudy (anonymous):

another way of looking at it is that \[\sin(630)=-1\] and \[\arcsin(-1)=-90\)

OpenStudy (zarkon):

if it is in radians (which is a little weird) then the answer is \(201\pi-630\)

OpenStudy (anonymous):

\[\arcsin(-1)=-90\] not sure where the \(\pi\) comes from

OpenStudy (anonymous):

how could it be in radians?

OpenStudy (zarkon):

I'm just going by what they said...one would think that it would be in degrees though

OpenStudy (anonymous):

if it is in radians then the answer would be \(-\frac{\pi}{2}\) or am i missing something?

OpenStudy (anonymous):

Zarkon, how did you get that answer?

OpenStudy (anonymous):

yeah enquiring minds want to know

OpenStudy (zarkon):

the 630 radians (which is not a multiple of pi)

OpenStudy (anonymous):

ooooh

OpenStudy (p0sitr0n):

just take 630 mod PI

OpenStudy (zarkon):

you need to keep subtracting \(2\pi\) to get back between \(-\pi/2\) and \(\pi/2\)

OpenStudy (anonymous):

more trickonometry

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

ohhh okayyy thanks all of you!!

OpenStudy (freckles):

\[\sin(630)=\sin(630-100*2\pi)\] This would mean we are now looking at 630-200pi which is in the second quadrant if we have are angle is between pi/2 and pi then we can say sin(pi-that angle)=sin(of some angle between 0 and pi/2) so \[\sin(630)=\sin(630-100*2\pi)=\sin(pi-[630-100 \cdot 2 \pi])\]

OpenStudy (anonymous):

Why did you choose to multiply 100 by \[2\Pi \]?

OpenStudy (freckles):

sin(x)=sin(x-k*2pi)

OpenStudy (freckles):

for any integer k

OpenStudy (freckles):

where x is in radians of course

OpenStudy (freckles):

if we had x in degrees then sin(x)=sin(x-k*360 deg)

OpenStudy (freckles):

k represents the number of rotations we make a round the circle

OpenStudy (anonymous):

Where did you get that equation from?

OpenStudy (freckles):

It is just something from going around and around again

OpenStudy (freckles):

but yeah it is a formula somewhere

OpenStudy (freckles):

it is a formula here

OpenStudy (anonymous):

ohh okay thanks!!!! :D

OpenStudy (freckles):

 say that is angle we are looking at right there